OpenStudy (anonymous):

-5 1/7e - (-5 2/5)e = -1/2

OpenStudy (anonymous):

I had to use slashes but the 1/7 2/5 an 1/2 are all fractions

OpenStudy (anonymous):

The left side evaluated is 9e/35, not -1/2

OpenStudy (anonymous):

My left side interpretation was:\[-\left(\left(5+\frac{1}{7}\right) e\right)--\left(5+\frac{2}{5}\right) e \]